Francisco leased equipment from Julio on December 31, 2021. The lease is a 10-year lease with annual payments of $160,000 due on December 31 of each year beginning December 31, 2021. The present value of the lease payments is $1,045,928. Francisco's incremental borrowing rate is 13% for this type of lease. The implicit rate of 11% is known by the lessee. What should be the balance in Francisco lease liability at December 31, 2022?

Multiple Choice

  • $819,880.

  • $885,928.

  • $829,890.

  • $823,380.
Answer #1
Lease liability at inception 1045928
Less: Payment on December 31, 2021 160000
Lease liability at December 31, 2021 885928
Interest for December 31,2022 97452 =885928*11%
Lease liability reduction for December 31,2022 62548 =160000-97452
Lease liability at December 31, 2022 823380 =885928-62548
Option D $823,380 is correct